OpenStudy (anonymous):

if a+b=45 then prove that (cota-1)(cotb-1)=2

OpenStudy (anonymous):

hy

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

did u know

OpenStudy (anonymous):

,cotacotb-1/cota+cotb=1 cota+cotb=cotacotb-1 cotacotb-cota-cotb=2.

OpenStudy (anonymous):

there's basically the answer fill in the rest by yourself.

OpenStudy (anonymous):

hy

OpenStudy (anonymous):

hy math

OpenStudy (loser66):

a+b =45 ---> a = b = 45-a \[cot a =\frac{1}{tan a}\] so we can re write like \[(\frac{1}{tana}-1)*(\frac{1}{tan(45-a)}-1)\] now, make the same denominator for the terms inside the bracket \[(\frac{1-tana}{tana})*(\frac{1-tan(45-a)}{tan(45-a)})\]

OpenStudy (loser66):

moreover, \[tan(45-a) = \frac{tan45-tana}{1+tan45tana}\] and tan45=1 replace into this we have\[ tan (45 -a) = \frac{1-tana}{1+tana}\] now plug the whole thing into the form we have

OpenStudy (loser66):

sorry , near answer but I have to go, will be back to finish, or someone else can continue my work.

OpenStudy (anonymous):

hy

OpenStudy (loser66):

you are the asker? ok, i am back, too late but if you still need it. I can continue.

OpenStudy (anonymous):

hy 66

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u can continue

OpenStudy (loser66):

I see many smarties here, let them contribute,ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Could you first show us what you've done with the question please?

OpenStudy (anonymous):

hy please help me

OpenStudy (anonymous):

@bijay56 Are you there? It's pointless for us to help you if you don't even attempt the question. It's for your benefit. We're not gaining anything if we just do the question for you.

OpenStudy (anonymous):

question-if a+b=45 then prove that (cota-1)(cotb-1)=2

OpenStudy (anonymous):

Could you show what you've done please?

OpenStudy (anonymous):

ans-cot(a+b)=cot45

OpenStudy (anonymous):

or,cotb.cota-1/cotb+cota=1

OpenStudy (anonymous):

Those are just the answers. I want you to show me your working out.

OpenStudy (anonymous):

or,cotb.cota-1=cotb+cota

OpenStudy (anonymous):

now what to do

OpenStudy (anonymous):

Have you attempted the question yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think u wont help me

OpenStudy (anonymous):

hy 66

OpenStudy (anonymous):

hy mertsj

OpenStudy (loser66):

Mertsj, why it take that long, you " are typing replies" in 11 minutes!!!???

OpenStudy (anonymous):

bijay56, if you've attempted the question, could we at least have a look at what you've done so far?

OpenStudy (loser66):

this is a new comer for sure. he/she even not know how to tag people. poor guy!!

OpenStudy (anonymous):

No point us helping you if you don't show us what you've done. Showing us what you've done is an important aspect of showing us that you want to learn and understand what we're trying to do for you.

OpenStudy (loser66):

@Azteck tolerate, he is off line, we come up with the problem on many ways, it's hard for the asker follow.

OpenStudy (anonymous):

Then we just have to wait until he/she comes back online. Then we can start helping the person.

OpenStudy (loser66):

you see, he said "hy 66" "hy 56" . he is poor.

OpenStudy (mertsj):

If a+b=45 then b = 45-a \[(\cot (a)-1)(\cot (45-a)-1)=2\] \[(\frac{1}{\tan (a)}-1)(\frac{1}{\tan (45-a)}-1)=2\] \[(\frac{1}{\tan a}-1)(\frac{1+\tan 45\tan a}{\tan 45-\tan a}-1)=2\] \[(\frac{1}{\tan a}-1)(\frac{1+\tan a}{1-\tan a}-1)=2\] \[(\frac{1-\tan a}{\tan a})(\frac{1+\tan a-(1-\tan a)}{1-\tan a})=2\] \[(\frac{1}{\tan a})(\frac{2\tan a}{1})=2\] \[2=2\]

OpenStudy (loser66):

and Mertsj, I guess you are kicked out the net. cannot take that long to reply

OpenStudy (loser66):

thanks Mertsj, don't delete, to the poor guy, definitely he will be back to take it.

OpenStudy (loser66):

can go to sleep now. bye bye everybody. hehehe

OpenStudy (anonymous):

a+b=45, =>cot(a+b)=cot45 =>(cota + cotb)/(cotacotb-1)=1 =>cota + cotb=cotacotb-1 =>cotacotb -cota -cotb -1=0 =>cotacotb -cota -cotb +1 = 2 =>(cota-1)(cotb-1)=2